OpenStudy (anonymous):

ON Feb. 10 Oscar rented a Chevy Trailblazer w/ a full tank of gas and having only 13091 miles on the odometer. On Feb. 12 he returned the vehicle w/ 13322 miles. The rental agency charged $92 for the rental and needed 14 gals of gas to fill the tank. a) find the SUVs rate of gas consumption in mpg b) find average cost of rental dollars per day c)find average rate of travel, miles per day d)find rental rate, in cents per mile

OpenStudy (ganpat):

a) 16.5 mpg b) 46 $ c) 115.5 miles per day d) 251 cents... assuming 10 - 12 Feb = 2 days...

OpenStudy (ganpat):

u want explanation for each ??

OpenStudy (anonymous):

yes please

OpenStudy (ganpat):

a) 13322 - 13092 = 231.. thats difference when car rented and when oscar gave back.. so 231/ 14 = 16.5 ... 14 = gal is amount of fuel used in tank..

OpenStudy (ganpat):

b) He rented it for 2 days at 92 $.. Therefore avg cost = 92/2 = 46 $

OpenStudy (anonymous):

a) find the SUVs rate of gas consumption in mpg we know the person drove 231 miles b/c 13322 - 13091 = 231 miles we also know the person used 14 galons of gas total so 231/14 = 16.5 mpg b) find average cost of rental dollars per day we know the person had the car for two days the rental agency charged $92 (assuming that was just the cost for the rental) 92/2 = 46 so the rental company charged $46 a day c) find average rate of travel, miles per day we know they traveled 231 miles and had the car for 2 days so 231/2 = 115.5 d) find rental rate, in cents per mile 251 cents

OpenStudy (ganpat):

c) average rate of travel = 231 miles in 2 days so for a day 231/2 = 115.5..

OpenStudy (ganpat):

d) rental rate cents per mile = total miles / total rent .. = 231/ 92 = 2.51 $ = 251 cents..

OpenStudy (ganpat):

gotta ?????

OpenStudy (anonymous):

I'm trying to figure out if there are equations i'm supposed to be turning these solutions into.

OpenStudy (ganpat):

well this are all basic average formula`s.. Need not to be equation..

OpenStudy (anonymous):

last few days we've been turning everything into equations . . . just checking

OpenStudy (ganpat):

k, let me know, how can i help u !!!